OpenStudy (anonymous):

What value of p will make this equation true? (6p + 4)/6 = (4p - 8)/3 f 2 g -6 h 10 j

OpenStudy (anonymous):

try getting rid of the denominators first...

OpenStudy (texaschic101):

(6p + 4) / 6 = (4p - 8) / 3 ---- multiply by the LCD, which is 6 6p + 4 = 2(4p - 8) 6p + 4 = 8p - 16 6p - 8p = -16 - 4 -2p = -20 p = 10 check.. 6p + 4 = 8p - 16 6(10) + 4 = 8(10) - 16 60 + 4 = 80 - 16 64 = 64 (correct) p = 10

OpenStudy (tkhunny):

(6p + 4)/6 = (4p - 8)/3 Multiply by 3 (6p + 4)/2 = (4p - 8) Factor 2(3p + 2)/2 = 4p - 8 Simplify 3p + 2 = 4p - 8 Such operations are easy enough to imagine. Just use the ones you need. The order is not that important.